OpenStudy (anonymous):

Please help again. :) 1/3 (6x-18) =16 Can you help me solve this please?

OpenStudy (anonymous):

1/3 (6x - 18) = 16 (6x - 18) = 16 * 3 = 48 6x = 48 + 18 6x = 66 x = 66/6 = 11

OpenStudy (anonymous):

expand ==> 2x+6 =16 ==> 2x=10 ===> x=5

OpenStudy (anonymous):

Let me check guys and thanks. You l get your metal when i figure who's right.

OpenStudy (anonymous):

*give

OpenStudy (anonymous):

Thanks Educatedgirl!

OpenStudy (anonymous):

Your welcome! :)

OpenStudy (anonymous):

I'll still give @reheen a medal. He/she at least tried. Thank you for that! :D